# Review of Demographic Fairness in Face Recognition

### v2

@@ II-B Variability in Skin-tone
The influence of skin-tone on the performance of FR systems has been extensively studied, revealing significant demographic disparities. In [31], Muthukumar et al. identified notable under-performance in recognizing dark-skinned females compared to other demographic group for commercial classifiers. Their analysis attributed these disparities to structural features such as lips, eyes, and cheeks, in addition to skin-tone itself. In another study, Krishnapriya et al.  [22] examined FMR and FNMR across skin-tone groups, observing higher FMR for African-American cohorts and higher FNMR for Caucasians; however, they did not find a direct causation between darker skin-tones and higher error rates. Buolamwini and Gebru [33] employed the Fitzpatrick skin classification system to evaluate commercial gender classifiers and reported the lowest accuracy for darker-skinned females. While this study evaluated gender classification rather than identity recognition, it led to broader discussions around fairness in facial analysis systems, including those used in biometric applications.
@@
The Biometric Technology Rallies organized by MdTF have offered comprehensive insights into the role of skin-related factors in FR bias. Their 2019 report [32] emphasized skin reflectance as a more significant predictor of performance disparities than race. Using systematic linear modeling, their study demonstrated that darker skin-tones were associated with longer transaction (processing overall pipeline) times and lower accuracy in biometric systems. The longer transaction times were primarily attributed to difficulties in the face detection or image acquisition stage under suboptimal lighting conditions. Lower skin reflectance can reduce contrast, making it harder for detection algorithms to localize the face, thus increasing processing time. This dependency was found to vary substantially across systems, highlighting important role of acquisition methods in determining the extent of bias. Lu et al.  [34] provided a quantitative assessment of performance variations across five skin-tone groups, identifying light-skinned individuals as the easiest to verify and darker-skinned individuals as the most challenging. However, ambiguities in defining skin-tone categories complicate direct evaluations, highlighting the need for standardized classification metrics.
@@
While many studies report that individuals with lighter skin-tones tend to be recognized more accurately than those with darker skin-tones, there is no consistent consensus that skin-tone is the primary driver of demographic bias. Moreover, it is important to note that most analyses rely on apparent or perceived skin tone–estimated via brightness or proxy classification schemes–rather than direct, objective measurement.
@@
TABLE I: Summary of works delving into various causes of demographic fairness in face recognition. As several works have identified multiple causes of bias, we have categorized the works based on their primary focus or inference. For details, readers are encouraged to refer to the source materials. The demographic factors of primary interest are denoted as ET: Ethnicity or race, GN: gender or sex, AG: age; whereas + indicates study of more attributes.
@@
Klare et al. [23] 2012 PCSO ET, GN, AG Advocated for balanced datasets and exclusive cohorts to improve face recognition performance.
@@
Muthukumar et al. [31] 2018 PPB GN, ST Identified structural facial features and skin tone as key factors for dark-skinned females’ underperformance.
@@
Phillips et al. [35] 2011 FRVT ET Identified the “other-race effect,” where algorithms performed better on their respective majority racial groups.
@@
Wang et al. [29] 2019 RFW ET Demonstrated racial bias using the RFW dataset, with higher error rates for African faces compared to Caucasians.
Serna et al. [37] 2019 DiveFace ET, GN Highlighted significant performance gaps across demographic groups, calling for fairness-aware algorithm designs.
@@
Lu et al. [34] 2019 IJB-B, IJB-C AG, GN, ST, + Analyzed multiple covariates; noted lighter skin tones consistently outperformed medium-dark tones.
@@
Ricanek et al. [39] 2015 ITWCC AG Discussed recognition challenges due to structural changes in children’s facial features over time.
@@
Albiero et al. [52] 2022 MORPH, MFAD Face region Studies impact of facial regions on bias patterns.
@@
II-C Algorithmic Factors
In this section, we examine the sensitivity and limitations of algorithms in addressing demographic attributes. Phillips et al. [35] identified the “other-race effect,” where algorithms developed in Western and East Asian contexts demonstrated superior performance for their respective majority racial groups. This disparity persisted even when datasets were balanced, pointing to underlying biases in algorithmic design and training processes. Klare et al.  [23] observed recognition challenges for specific demographic groups, including females, Black individuals, and younger cohorts. They reported improved performance when models were trained exclusively on specific demographic cohorts. This improvement was supported by the integration of dynamic face matchers within the recognition pipeline—a modification that reflects both algorithmic adaptation and the influence of training cohort composition. In [36], Nagpal et al. demonstrated that deep learning models encode in-group biases, mirroring human tendencies such as own-race and own-age effects. By analyzing activation maps, they showed that these biases were ingrained within the feature representations of the models. Wang et al. [29], using the Racial Faces in-the-Wild (RFW) dataset, validated racial bias, revealing that error rates for African faces were nearly double those for Caucasians, even with race-balanced training data. Serna et al. [37] similarly highlighted significant performance gaps across demographic groups, advocating for diverse training datasets and fairness-aware algorithmic designs.
@@ II-C Algorithmic Factors
Further investigations into gender-based disparities by Albiero et al. [38] revealed skewed impostor and genuine score distributions as the primary reasons for lower accuracy in women. This disparity persisted across datasets, regardless of balanced training and neutral facial expressions. Ricanek et al. [39] noted unique challenges in recognizing children’s faces due to structural changes with age, finding that algorithms effective on adult faces performed poorly for younger subjects. Although this issue pertains to longitudinal variations in facial morphology, it underscores a broader implication: the same algorithms may not be equally effective across seemingly similar tasks (FR). Attributes such as facial structure should therefore be considered as critical factors in guiding the design and selection of appropriate recognition algorithms.
@@
Identifying algorithmic factors that contribute to differential recognition accuracy across demographic groups remains a complex task. This process requires isolating the influence of the training algorithm–an integral component of the recognition pipeline. To ensure valid attribution, experimental setups need to be carefully controlled that differ in nothing but algorithms. Initially, performance disparities were primarily attributed to handful factors such as imbalances in training datasets or variations in acquisition conditions. However, as the number of potential influencing factors grows, the experimental studies to rigorously investigate and isolate the algorithmic sources of bias require a thorough and systematic approach.
II-D Image Quality
@@
The quality of input images and associated covariates significantly influence the manifestation of demographic differentials in FR systems. Numerous studies have emphasized how disparities in image quality across different demographic groups can lead to variations in system performance. For instance, Cavazos et al. [24] analyzed both data-driven and scenario-based factors, revealing that dataset complexity and decision thresholds have a notable impact on recognition accuracy and racial bias. Their experiments across multiple algorithms further demonstrated that East Asian faces required higher decision thresholds compared to Caucasian faces to achieve equivalent error rates, highlighting the interplay between dataset characteristics and demographic attributes.
@@
II-E Combined or Intersectional Factors
@@
Age-related biases have been linked to structural transformations in facial features over time, particularly among children. Ricanek et al. [39] observed increased complexity of child aging compared to adults, attributing recognition challenges to changes in facial bone structure and the proportions of facial components. Additionally, Best-Rowden and Jain [47] reported nuanced patterns in age-related recognition performance, noting that while males generally exhibit higher genuine scores, their performance declines more rapidly with age compared to females. These observations underline the intricate interplay of demographic attributes in shaping biases in FR systems. The intersection of age, race, and gender significantly amplifies accuracy differentials in FR systems. In [48], Sarridis et al. identified a disproportionately high number of incorrect verifications (both false positives and false negative)for African females over 60 years compared to Caucasians, illustrating the compounded effects of intersecting demographic factors. Similarly, El Khiyari et al. [49] demonstrated that face verification accuracy is notably lower for younger individuals (aged 18–30), females, and certain racial groups such as Black individuals, highlighting the challenges posed by such intersections of demographic factors. Liu et al.  [62] observed that multiple factors such as ethnicity, pose, occlusion, and blur jointly impact FR accuracy.
Among the few studies addressing explanatory frameworks, [50] introduced a methodology aimed at evaluating the influence of both protected and non-protected attributes on FR performance. Their analysis revealed that fairness outcomes exhibit considerable variability across multi-attribute groupings. Notably, patterns observed in single-attribute evaluations frequently diminish or invert when extended to analyses involving multiple attributes.
@@
The inconsistent findings indicate that age-related bias in FR remains insufficiently understood and lacks a clear consensus. Overall these studies highlight the complexity of mitigating intersecting biases within FR systems, underscoring the importance of incorporating age-specific and multi-factor considerations in both algorithm development and evaluation protocols.
II-F Soft Attributes or Non-Demographic Factors
@@
III Datasets for the Study of Demographic Accuracy Differences
In evaluating and mitigating demographic fairness in FR systems, the selection of suitable datasets plays an important role. Although numerous FR datasets exist, those specifically intended for bias and fairness-related tasks must include demographic labels associated with each subject or identity. The datasets designed for tasks such as race, gender, ethnicity, or age estimation are particularly useful when they include demographic labels, as such factors are critical for assessing fairness. For certain tasks (related to estimation or classification of attributes), having a single image per subject may suffice, as training and testing can be conducted separately for each image. However, FR models, especially state-of-the-art systems, benefit from having multiple images (variations) of each identity to train more robust feature extractors. The testing phase or deployment has two primary modes: one-to-one (1:1) verification and one-to-many (1:$N$) identification. In 1:1 matching, a probe image is compared against one or more enrolled images of a claimed identity. For identification (1:$N$ matching), the probe is compared against a gallery containing reference images from multiple identities to determine the closest match. Occasionally, the gallery may be composed of multiple images per identity.
TABLE III: Commonly used datasets for tasks related to demographic accuracy differentials in face recognition. Appendix VII provides the URLs of the referenced datasets.
@@ III Datasets for the Study of Demographic Accuracy Differences
BFW  [76] [77] 2023 20,000 / 800 (Male, Female), (Asian, Black, Indian, White) Test
@@
CausalFace  [79] 2023 48,000 / 10,200 (Male, Female), (Asian, Black, White) Test
These requirements significantly reduce the availability of datasets suitable for assessing demographic fairness, as most FR datasets do not provide adequate demographic labels or have highly skewed distributions of subjects across different demographic groups. It is important to note that several prominent bias evaluation efforts, such as those led by NIST and the MdTF, rely on proprietary or restricted-access datasets. While these organizations have conducted large-scale, operationally relevant evaluations, their limited availability to the wider research community poses challenges for reproducibility, benchmarking, and comparative analysis. In this section, we outline recent and commonly used publicly available datasets that are relevant for assessing and addressing demographic fairness in FR systems.
@@
  • VGGFace2 [71]: The VGGFace2 dataset is a large-scale FR dataset containing over 3.31 million images of 9,131 subjects. This dataset was annotated for gender (Male, Female) and ethnicity (Asian, Black, Indian, White) labels by Idiap Research Institute [^1] Annotations for VGGFace2 Dataset making it useful for fairness related tasks. This dataset offers significant advantages due to a large number of identities with variations in age, pose, and lighting; VGGFace2 improves over the original with more diversity too. The dataset, however, is significantly imbalanced, with a disproportionate representation of White and male subjects. Additionally, the original download source has been removed by the creators as of 2024.
  • DemogPairs [72]: The DemogPairs dataset is a validation set containing 10.8K images, divided into six demographic folds: Asian females, Asian males, Black females, Black males, White females, and White males. The dataset was specifically designed to evaluate the demographic bias in FR models, offering 58.3 million evaluation pairs, including cross-demographic, cross-gender, and cross-ethnicity pairs. The DemogPairs dataset was constructed with rigorous demographic annotation and is a useful resource for testing the generalization of FR systems across diverse demographic groups.
@@
  • CausalFace [79]: It is a large-scale dataset of synthetically generated faces comprising 48,000 synthetic face image pairs generated from 10,200 unique identities, along with 555,000 human annotations covering individual attributes and pairwise identity comparisons. The dataset is constructed using EG3D [84], a state-of-the-art GAN framework that allows explicit control over geometry and pose. In terms of demographic attributes, the CausalFace dataset includes six demographic groups created by combining three ethnicities (White, Black, and East Asian) and two genders (male and female). Using a GAN-based generator, face images were systematically modified across unprotected attributes such as pose, age, expression, and lighting while maintaining identity consistency. The dataset images were selected on the basis of perceived matching scores provided by human annotators. Although it offers a bias-free ground truth, its synthetic nature may limit generalizability to real-world scenarios.
The datasets are summarized in Table III, and Fig. 2 illustrates an overview of the demographic distribution, with race and gender as demographic variables, for the datasets reviewed in this section.
IV Assessment of Demographic Fairness in FR
For an FR system, the FMR represents the proportion of matching scores where a non-mated pair of biometric samples (i.e., samples from different subjects) is incorrectly classified as a match, while FNMR corresponds to the scenarios where a mated pair (i.e., samples of the same subject) is incorrectly classified as non-matching. These metrics are essential for evaluating the performance of FR algorithms– where lower values of error rates are preferred. These errors are determined by the score threshold (often denoted as $\tau$) that binarizes the similarity score into match or no-match decision. As illustrated in Figs. 1(a)1(c), the overlap between distributions of mated scores (shown in green) and non-mated scores (shown in red) defines the regions where classification errors (i.e., false matches and false non-matches) occur.
The choice of score threshold plays a significant role in FMR and FNMR. A higher threshold reduces FMR by limiting the likelihood of imposter pairs being classified as matches but increases FNMR by rejecting more genuine pairs, and vice-versa. However, it is essential to distinguish between score overlap and binary decision-making. The overlap reflects the inherent ambiguity in the score distributions, while binarized decisions result from applying the threshold to assign matches or non-matches. For instance, two demographic groups with identical overlap may exhibit different FMR and FNMR due to variations in their score distribution shapes or population characteristics. Similarly, the distributions (mated, non-mated, or both) for two demographic groups can be significantly different albeit exhibiting same error rates.
The demographic bias in FR system leads to variations in score distributions and their overlaps across different demographic groups. Such disparities inherently result in different FMR and FNMR values for each group when a single (global) threshold is used. Fig. 3 illustrates score distributions of different groups may differ, leading to unequal error rates. In this section we briefly review performance measures designed for demographic-aware assessments of fairness in FR systems. Given the overlap and distinct characteristics of scores and decisions, establishing a well-defined evaluation framework is crucial for accurate analysis. Howard et al.  [6] introduced the concepts of differential performance and differential outcomes, providing two key terms that aid in achieving a precise understanding and categorization of assessment metrics.

### v3

@@ III-B Variability in Skin-tone
In [36], Muthukumar et al. identified notable under-performance in recognizing dark-skinned females compared to other demographic group for commercial classifiers. Their analysis attributed these disparities to structural features such as lips, eyes, and cheeks, in addition to skin-tone itself. In another study, Krishnapriya et al.  [26] analyzed FMR and FNMR across skin-tone groups and observed error rate variations among demographic cohorts; however, they did not find conclusive evidence linking darker skin-tones directly to higher error rates. Buolamwini and Gebru [40] employed the Fitzpatrick skin classification system to evaluate commercial gender classifiers and reported the lowest accuracy for darker-skinned females. While this study evaluated gender classification rather than identity recognition, it led to broader discussions around fairness in facial analysis systems, including those used in biometric applications.
@@
The Biometric Technology Rallies organized by MdTF have offered comprehensive insights into the role of skin-related factors impacting FR performance. Their 2019 report [37] emphasized skin reflectance as a more significant predictor of performance disparities than race. Skin reflectance refers to the measurable amount of light reflected from the skin surface, which affects how well facial features are captured in images; whereas skin-tone refers to perceived skin color [37]. Using systematic linear modeling, their study demonstrated that darker skin-tones were associated with longer transaction (processing overall pipeline) times and lower accuracy in biometric systems. The longer transaction times were primarily attributed to difficulties in the face detection or image acquisition stage under suboptimal lighting conditions. Lower skin reflectance can reduce contrast, making it harder for detection algorithms to localize the face, thus increasing processing time. This dependency was found to vary substantially across systems, highlighting important role of acquisition methods in determining the extent of performance differences. Lu et al.  [41] provided a quantitative assessment of performance variations across five skin-tone groups, identifying light-skinned individuals as the easiest to verify and darker-skinned individuals as the most challenging. However, ambiguities in defining skin-tone categories complicate direct evaluations, highlighting the need for standardized classification metrics.
@@
While many studies report that individuals with lighter skin-tones tend to be recognized more accurately than those with darker skin-tones, there is no consistent consensus that skin-tone is the primary driver of differences in FR performance across demographic groups. Moreover, it is important to note that most analyses rely on apparent or perceived skin tone–estimated via brightness or proxy classification schemes–rather than direct, objective measurement.
@@
TABLE I: Summary of works delving into various causes of demographic fairness in face recognition. As several works have identified multiple causes of such performance differences, we have categorized the works based on their primary focus or inference. For details, readers are encouraged to refer to the source materials. The demographic factors of primary interest are denoted as ET: Ethnicity or race, GN: gender or sex, AG: age; whereas + indicates study of more attributes.
@@
Training Datasets Klare et al. [27] 2012 PCSO ET, GN, AG Advocated for balanced datasets and exclusive cohorts to improve face recognition performance.
@@
Skin-Tone Muthukumar et al. [36] 2018 PPB GN, ST Identified structural facial features and skin tone as key factors for dark-skinned females’ underperformance.
@@
Algorithmic Factors Phillips et al. [42] 2011 FRVT ET Identified the “other-race effect,” where algorithms performed better on their respective majority racial groups.
@@
Dooley et al. [46] 2023 CelebA, VGGFace2, RFW GN Demonstrated that network architecture and hyperparameters inherently influence fairness in FR; and bias can arise from model inductive biases.
Image Quality Lu et al. [41] 2019 IJB-B, IJB-C AG, GN, ST, + Analyzed multiple covariates; noted lighter skin tones consistently outperformed medium-dark tones.
@@
Intersectional Factors Ricanek et al. [43] 2015 ITWCC AG Discussed recognition challenges due to structural changes in children’s facial features over time.
@@
Soft-Biometric Attr. Albiero et al. [59] 2022 MORPH, MFAD Face region Studies impact of facial regions on bias patterns.
@@
III-C Algorithmic or Model Factors
While majority of sources of performance differences in FR can be traced back to data, several studies have identified how specific algorithm-related choices can amplify or introduce demographic disparities. These algorithmic elements often interact with other factors/ attributes, making causal attribution challenging. In this section, we examine works where algorithm-level factors are prominently discussed as contributors to observed bias. Phillips et al. [42] identified the “other-race effect,” where algorithms developed in Western and East Asian contexts demonstrated superior performance for their respective majority racial groups. The said effect refers to an observation that humans recognize faces of their own race more accurately than faces of other races. This disparity persisted even when datasets were balanced, hypothesizing that underlying biases may occur due to algorithmic design and training processes. However, it may be noted that most of these findings are based on systems developed over a decade ago, and their implications may vary for modern deep learning-based FR models. Klare et al.  [27] observed recognition challenges for specific demographic groups, including females, Black individuals, and younger cohorts. They reported improved performance when models were trained exclusively on specific demographic cohorts. This improvement was supported by the integration of dynamic face matchers within the recognition pipeline–a modification that reflects both algorithmic adaptation and the influence of training cohort composition. Similarly, several recent works proposed modifying features or decision thresholds based on group information (see Sec. VI). While such approaches might lead to improved accuracy for under-represented groups, it requires reliable access to demographic labels. This introduces few drawbacks: external labeling increases operational cost and may be prohibited under privacy or anti-discrimination regulations (e.g., GDPR). Alternatively, predicting demographic attributes automatically introduces another layer of uncertainty, where misclassification can directly impact system fairness and performance.
In [44], Nagpal et al. demonstrated that deep learning models encode in-group biases, mirroring human tendencies such as own-race and own-age effects. By analyzing activation maps, they showed that these biases were ingrained within the feature representations of the models. Based on these observations, they hypothesized that classification models encode faces from different racial groups in distinct ways, indicating the presence of bias. Although the study did not isolate algorithmic design as the sole cause, it highlighted how internal model behavior may reflect or reinforce demographic disparities. Dooley et al. systematically evaluated the impact of network architectures and hyper-parameters on fairness in FR [46]. Their findings revealed that architectural choices significantly influence demographic disparities. It also demonstrated that bias can be inherently encoded in a model’s inductive bias, and relying solely on standard architectures optimized for accuracy may compromise fairness.
@@ III-C Algorithmic or Model Factors
Further investigations into gender-based disparities by Albiero et al. [45] revealed skewed impostor and genuine score distributions resulting in a lower accuracy in women. This disparity persisted across datasets, regardless of balanced training and neutral facial expressions, and hence, the authors speculated that the cause could be more intrinsic factors between the groups. Ricanek et al. [43] noted unique challenges in recognizing children’s faces due to structural changes with age, finding that algorithms effective on adult faces performed poorly for younger subjects. Although this issue pertains to longitudinal variations in facial morphology, it underscores a broader implication: the same algorithms may not be equally effective across seemingly similar tasks (FR). Attributes such as facial structure should therefore be considered as critical factors in guiding the design and selection of appropriate recognition algorithms.
@@
Identifying algorithmic factors that contribute to differential recognition accuracy across demographic groups remains a complex task. This process requires isolating the influence of the training algorithm–an integral component of the recognition pipeline. To ensure valid attribution, experimental setups need to be carefully controlled that differ in nothing but algorithms. Initially, performance disparities were primarily attributed to handful factors such as imbalances in training datasets or variations in acquisition conditions. However, as the number of potential influencing factors grows, the experimental studies to rigorously investigate and isolate the algorithmic sources of performance differences require a thorough and systematic approach.
III-D Image Quality
@@
The quality of input images and associated covariates significantly influence the manifestation of demographic differentials in FR systems. Numerous studies have emphasized how disparities in image quality across different demographic groups can lead to variations in system performance. For instance, Cavazos et al. [28] analyzed both data-driven– focusing on aspects of image quality– and scenario-based factors, revealing that dataset complexity and decision thresholds have a notable impact on recognition accuracy and racial bias. Their experiments across multiple algorithms further demonstrated that East-Asian faces required higher decision thresholds compared to Caucasian faces to achieve equivalent error rates. The authors also acknowledged that image quality may not act in isolation and that multiple underlying causes may jointly contribute to observed disparities.
@@
III-E Combined or Intersectional Factors
@@
Age-related disparities in FR performance have been linked to structural transformations in facial features over time, particularly among children. Ricanek et al. [43] observed increased complexity of child aging compared to adults, attributing recognition challenges to changes in facial bone structure and the proportions of facial components. Additionally, Best-Rowden and Jain [54] reported nuanced patterns in age-related recognition performance, noting that while males generally exhibit higher genuine scores, their performance declines more rapidly with age compared to females. These observations underline the intricate interplay of demographic attributes in shaping performance of FR systems. The intersection of age, race, and gender significantly amplifies accuracy differentials in FR systems. In [58], Sarridis et al. identified a disproportionately high number of incorrect verifications (both false positives and false negative)for African females over 60 years compared to Caucasians, illustrating the compounded effects of intersecting demographic factors. Similarly, El Khiyari et al. [53] demonstrated that face verification accuracy is notably lower for younger individuals (aged 18–30), females, and certain racial groups such as Black individuals, highlighting the challenges posed by such intersections of demographic factors. Liu et al.  [69] observed that multiple factors such as ethnicity, pose, occlusion, and blur jointly impact FR accuracy.
Among the few studies addressing explanatory frameworks, [57] introduced a methodology aimed at evaluating the influence of both protected and non-protected attributes on FR performance. Here, protected attributes refer to human characteristics legally safeguarded from discrimination (e.g., race or gender), whereas non-protected attributes include modifiable traits such as makeup or hairstyle. Their analysis revealed that fairness outcomes exhibit considerable variability across multi-attribute groupings. Notably, patterns observed in single-attribute evaluations frequently diminish or invert when extended to analyses involving multiple attributes.
@@
The inconsistent findings indicate that age-related performance differences in FR remains insufficiently understood and lacks a clear consensus. Overall these studies highlight the complexity of mitigating intersecting biases within FR systems, underscoring the importance of incorporating age-specific and multi-factor considerations in both algorithm development and evaluation protocols.
III-F Soft-Biometric Attributes or Non-Demographic Factors
@@
IV Datasets for the Study of Demographic Accuracy Differences
In evaluating and mitigating demographic fairness in FR systems, the selection of suitable datasets plays an important role. Although numerous FR datasets exist, those specifically intended for differential performance and fairness-related tasks must include demographic labels associated with each subject or identity. The datasets designed for tasks such as race, gender, ethnicity, or age estimation are particularly useful when they include demographic labels, as such factors are critical for assessing fairness. For certain tasks (related to estimation or classification of attributes), having a single image per subject may suffice, as training and testing can be conducted separately for each image. However, FR models, especially state-of-the-art systems, benefit from having multiple images (variations) of each identity to train more robust feature extractors. The testing phase or deployment has two primary modes: one-to-one (1:1) verification and one-to-many (1:$N$) identification. In 1:1 matching, a probe image is compared against one or more enrolled images of a claimed identity. For identification (1:$N$ matching), the probe is compared against a gallery containing reference images from multiple identities to determine the closest match. Occasionally, the gallery may be composed of multiple images per identity.
TABLE III: Commonly used datasets for tasks related to demographic accuracy differentials in face recognition.
@@ IV Datasets for the Study of Demographic Accuracy Differences
BUPT-GlobalFace [83] 2020 2M+ / 38,000 (Asian, African, Caucasian, Indian) Train/ Test
@@
BFW [85] [86] 2020 20,000 / 800 (Male, Female), (Asian, Black, Indian, White) Test
@@
These requirements significantly reduce the availability of datasets suitable for assessing demographic fairness, as most FR datasets do not provide adequate demographic labels or have highly skewed distributions of subjects across different demographic groups. It is important to note that several prominent bias evaluation efforts, such as those led by NIST and the MdTF, rely on proprietary or restricted-access datasets. While these organizations have conducted large-scale, operationally relevant evaluations, their limited availability to the wider research community poses challenges for reproducibility, benchmarking, and comparative analysis. In this section, we outline recent and commonly used publicly available datasets that are relevant for assessing and addressing demographic fairness in FR systems. The datasets presented below are arranged in chronological order of their release.
@@
  • Adience [77]: The Adience dataset consists of 26,580 images from 2,284 individuals captured in unconstrained (in the wild) conditions using typical consumer devices. The dataset is annotated with identity, gender, and one of eight age group labels. While primarily designed for age and gender classification tasks, the dataset’s natural variation in pose, illumination, and demographics makes it valuable for FR research. Sourced from nearly 200 publicly available Flickr albums, this dataset has a higher representation of female subjects.
@@
  • VGGFace2 [80]: The VGGFace2 dataset is a large-scale FR dataset containing over 3.31 million images of 9,131 subjects. This dataset was annotated for gender (Male, Female) and ethnicity (Asian, Black, Indian, White) labels by Idiap Research Institute [^1] Annotations for VGGFace2 Dataset making it useful for fairness related tasks. This dataset offers significant advantages due to a large number of identities with variations in age, pose, and lighting; VGGFace2 improves over the original with more diversity too. The dataset, however, is significantly imbalanced, with a disproportionate representation of White and male subjects. Although this dataset offers high intra-class variation with over 3 million images, it contains only slightly more than 9,000 unique identities, which may limit its utility for evaluating false positive rates in large-scale evaluations. Note that, the original download source has been removed by the creators as of 2024.
  • DemogPairs [81]: The DemogPairs dataset is a validation set containing 10.8K images, divided into six demographic folds: Asian females, Asian males, Black females, Black males, White females, and White males. The dataset was specifically designed to evaluate the demographic disparities in FR models, offering 58.3 million evaluation pairs, including cross-demographic, cross-gender, and cross-ethnicity pairs. The DemogPairs dataset was constructed with rigorous demographic annotation and is a useful resource for testing the generalization of FR systems across diverse demographic groups.
@@
  • KANFace [82]: The KANFace dataset comprises around 40,000 still images and 44,000 video clips (totaling roughly 14.5 million frames) of 1,045 individuals, collected in unconstrained settings. On average, each subject is represented by 39 images and over 13,000 video frames. The dataset includes manual annotations for identity, age, gender, and kinship, with demographic metadata sourced from Wikipedia and IMDb. The dataset consists of 586 male and 459 female subjects.
@@
  • BUPT-GlobalFace [83]: In addition to the BUPT-BalancedFace dataset, the BUPT-GlobalFace dataset is also widely used in fairness tasks. Unlike its balanced counterpart, BUPT-GlobalFace is racially imbalanced, containing over 2 million images from approximately 38,000 celebrities, with an over-representation of certain racial groups (e.g., Caucasian and Asian). Although it reflects real-world distribution more closely, the imbalance makes it less suitable for bias mitigation training and more appropriate for evaluating bias under unconstrained, skewed demographic settings.
@@
The datasets are summarized in Table III, and Fig. 3 illustrates an overview of the demographic distribution, with race and gender as demographic variables (where details of both variables were available), for the datasets reviewed in this section.
V Assessment of Demographic Fairness in FR
The bias in FR system leads to variations in score distributions and their overlaps across different demographic groups. Such disparities inherently result in different FMR and FNMR values for each group when a single (global) threshold is used. Fig. 1 illustrates score distributions of different groups may differ, leading to unequal error rates. In this section we briefly review performance measures designed for demographic-aware assessments of fairness in FR systems. Given the overlap and distinct characteristics of scores and decisions, establishing a well-defined evaluation framework is crucial for accurate analysis. Howard et al.  [10] introduced the concepts of differential performance and differential outcomes, providing two key terms that aid in achieving a precise understanding and categorization of assessment metrics.